OpenStudy (anonymous):

Expand the expression. log(base)3 sqrt. 9x Please show work. Thank you.

OpenStudy (anonymous):

\[\log_{3} \sqrt{9x}\]

OpenStudy (y2o2):

\[\log_{3}(\sqrt{9}) + \log_{3}(\sqrt{x}) = \log_{3}(3) +\log_{3}(\sqrt{x}) = 1+\log_{3}(\sqrt{x})\]

OpenStudy (anonymous):

Thank you soo much^_^ Now i'm understanding this a little!